OpenStudy (anonymous):

A solution of salt and water contains 60 grams of water per 120 milliliters of the solution. If 1 mole of water weighs 16 grams, _____ moles of water would be present in 40 milliliters of the solution. [Round off your answer to two decimal places] Numerical Answers Expected! @TheRealMeeeee @bibby @Dominick

OpenStudy (therealmeeeee):

well 60/16 is the moles and 40/120 is a third, i would say the answer of 60/16 mult by a third.

OpenStudy (therealmeeeee):

these are ratio problems kind of like this 1mole of water/16 g= x mole of water/60 g = 60/16 = moles of water then that answer/120 ml = how many moles/40 = 40(that answer)/120= how many moles in 40

OpenStudy (therealmeeeee):

16 x 3.75 = 60 so 3.75 moles

OpenStudy (anonymous):

1g H20 in 2mL soln => 20 g H2O in 40mL soln => 20g/16g/mol Answer: 1.25 moles

OpenStudy (therealmeeeee):

do u get it

OpenStudy (therealmeeeee):

16 x 3.75 = 60 so 3.75 moles

OpenStudy (therealmeeeee):

no punch in 40 times 3.75 and then divide by 120 you should get 1.25

OpenStudy (anonymous):

oh i got 2.75

OpenStudy (anonymous):

i multiplied wrong then

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

ok i got 3.75 now

OpenStudy (anonymous):

ok i got 1.25